OpenStudy (help!!!!):

Help me solve 3(2x-1)^2 +1=31

OpenStudy (anonymous):

so i would suggest moving the 1 to the other side and dividing everything by 3 meaning : ----> 3 (2x-1)^2 = 31-1 so 3(2x-1)^2 = 30 so after dividing everything by 3: (2x-1)^2 = 10 now simply open the brackets and move 10 to the other side to form a quadratic equation and solve